In a study of the accuracy of fast food drive through orders, one restaurant had 29 orders that were not accurate among 298 orders observed. Use a 0.05 significance level to test the claim that the rate of inaccurate orders is equal to 10%. Does the accuracy rate appears to be acceptable?
Answer #1

Solution :

This is the two tailed test .

The null and alternative hypothesis is

H0 : p = 0.10

Ha : p 0.10

= x / n = 29 / 298 = 0.0973

P0 = 0.10

1 - P0 = 0.90

Test statistic = z

= - P0 / [P0 * (1 - P0 ) / n]

= 0.0973 - 0.10 / [(0.10 * 0.90) / 298]

= -0.154

P(z < -0.154) = 0.4388

P-value = 2 * 0.4388 = 0.8776

= 0.05

P-value >

Fail to reject the null hypothesis .

The accuracy rate does not appears to be acceptable .